Arrow Distributing Corp.lkes to track inventory by using weeks of supply as well as by inventoryturnover. Arrow Distributing Corp Net Revenue Cost of sales Inventory Total assets $17,300 $12,280 $1,100 $9,350 a) What is its weeks of supply?weeks (round your response to two decimal placos)
Answer #1

a. Invetory Turn over= Cost of goods sold/Avg inventory

= 12280/1100=11.16

Weeks of supply= 52/ Inventory Turn over= 52/11.16= 4.66 weeks